OpenStudy (anonymous):

An airplane flew 3.5 hours with a 40 mph head wind. The return trip with a tail wind of the same speed took 2.5 hours. Find the speed of the plane in still air. 240 mph 220 mph 200 mph 180 mph

OpenStudy (anonymous):

Because the plane is traveling an equal distance both ways, you'll want to solve for x as follows: \[3.5(x-40)=2.5(x+40)\]\[3.5x-140=2.5x+100\]\[x=240\] The answer is 240 mph.

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

Yes, correct!